# TokenRec: Learning to Tokenize ID for LLM-based Generative Recommendations

### v1

@@ Abstract.
There is a growing interest in utilizing large-scale language models (LLMs) to advance next-generation Recommender Systems (RecSys), driven by their outstanding language understanding and in-context learning capabilities. In this scenario, tokenizing (i.e., indexing) users and items becomes essential for ensuring a seamless alignment of LLMs with recommendations. While several studies have made progress in representing users and items through textual contents or latent representations, challenges remain in efficiently capturing high-order collaborative knowledge into discrete tokens that are compatible with LLMs. Additionally, the majority of existing tokenization approaches often face difficulties in generalizing effectively to new/unseen users or items that were not in the training corpus. To address these challenges, we propose a novel framework called TokenRec, which introduces not only an effective ID tokenization strategy but also an efficient retrieval paradigm for LLM-based recommendations. Specifically, our tokenization strategy, Masked Vector-Quantized (MQ) Tokenizer, involves quantizing the masked user/item representations learned from collaborative filtering into discrete tokens, thus achieving a smooth incorporation of high-order collaborative knowledge and a generalizable tokenization of users and items for LLM-based RecSys. Meanwhile, our generative retrieval paradigm is designed to efficiently recommend top-$K$ items for users to eliminate the need for the time-consuming auto-regressive decoding and beam search processes used by LLMs, thus significantly reducing inference time. Comprehensive experiments validate the effectiveness of the proposed methods, demonstrating that TokenRec outperforms competitive benchmarks, including both traditional recommender systems and emerging LLM-based recommender systems.
@@
Recommender Systems, Large Language Models (LLMs), ID Tokenization, Vector Quantization, Collaborative Filtering (CF).
* Corresponding author: Wenqi Fan, Department of Computing, and Department of Management and Marketing, The Hong Kong Polytechnic University.
journalyear: 2024ccs: Information systems Recommender systems
1. Introduction
@@
As a prominent branch in the data mining field, recommender systems (RecSys) serve as an indispensable and effective technique in addressing information overload problems and enriching user experience across diverse applications (Chen et al., 2023; Zhao et al., 2024), such as e-commerce, job search, and social media platforms. To provide personalized recommendations that accord with user preferences, one of the most representative techniques is collaborative filtering (CF), which aims to capture collaborative knowledge by modeling the history of user-item interactions (Fan et al., 2019b). For example, as the most classic CF method, Matrix Factorization (MF) (Koren et al., 2009) decomposes the use-item matrix into two low-rank matrices (i.e., user&item representations) to represent each user and item with a unique ID (namely ID indexing/tokenizing), and calculates matching scores to predict user behaviors and generate recommendations via the inner product between their representations. Due to the superior ability in representation learning on graph-structured data, Graph Neural Networks (GNNs) (e.g., LightGCN (He et al., 2020) and GTN (Fan et al., 2022)) have been recently employed to significantly enhance CF by capturing higher-order collaborative knowledge on user-item interaction graphs for recommendations in IDs manners. The main idea of the existing methods is to obtain learnable representations (i.e., token embeddings) for discrete user&item IDs (i.e., tokens).
@@
Figure 1. Comparison of ID tokenization methods in LLM-based recommendations. Unlike the existing methods, our approach can tokenize users and items with LLM-compatible tokens by leveraging high-order collaborative knowledge.
@@
More recently, the rapid development of Large Language Models (LLMs) techniques (e.g., ChatGPT and LLaMA) has showcased notable milestones for revolutionizing natural language processing techniques (Brown et al., 2020; Zhao et al., 2023). Technically, the unprecedented power of LLMs can be attributed to the scaling up of model/parameter size alongside a tremendous amount of training corpus. In particular, LLMs equipped with billion-scale parameters have exhibited unprecedented language understanding and generation abilities, along with remarkable generalization capability and in-context learning skills that facilitate LLMs to better generalize to unseen tasks and domains (Zhao et al., 2023). Given the emerging trends and aforementioned advancements of LLMs, LLM-empowered recommender systems have drawn increasing attention from recent studies and demonstrated distinctive abilities for advancing recommender systems (Zhao et al., 2024). For example, ChatRec (Gao et al., 2023) integrates ChatGPT into conversational recommendations, which enables users to deliver their explicit preferences in natural language. P5 (Geng et al., 2022) introduces an LLM-based recommendation framework that unifies diverse recommendation tasks by multi-task prompt-based pre-training.
@@
Unlike the ID-based recommendation methods (e.g., MF and GNN-based), users&items tokenization is one of the most critical steps to take advantage of LLMs within recommendations, as shown in Figure 1. To be more specific, the naive approach known as Independent Indexing (IID) assigns special tokens (i.e., ID) to tokenize each user and item within language models directly. However, this approach becomes infeasible and unrealistic when dealing with large-scale real-world recommender systems, in which the sizes of users and items typically reach billions, dramatically expanding the token vocabulary in LLMs. Meanwhile, as a natural solution, textual title indexing is proposed to utilize LLMs’ in-vocabulary tokens to represent items based on their titles and descriptions, such as the example ”Apple iPhone 15, 256 GB, black”, thus avoiding vocabulary size explosion in LLMs (Gao et al., 2023; Bao et al., 2023). To achieve a closer alignment between recommendations and natural language, P5 employs whole-word embedding (Takase and Kobayashi, 2020) to indicate whether consecutive sub-word tokens originate from the same entities (i.e., user/item). Some studies use continuous embedding (i.e., soft indexing) learned from encoders to represent users&items in LLMs-based recommendations (Liao et al., 2023; Zhang et al., 2023).
Despite the success mentioned above, the majority of existing methods for users&items tokenization for LLM-based recommender systems still have several limitations. For example, the use of whole-word embedding cannot effectively capture high-order collaborative knowledge and generalize well to unseen users/items for recommendation. In addition, due to the nature of discrete tokens in language models, using continuous indexing makes it challenging to align LLMs in recommender systems closely.
@@
The remainder of this paper is structured as follows. Section 2 introduces the proposed approach, which is evaluated in Section 3. Then, Section 4 summarizes the recent development of collaborative filtering and LLM-based RecSys. Finally, conclusions are drawn in Section 5.
2. The Proposed Method
@@ 2.1. Notations and Definitions
(1) $\mathbf{z}_{i}={\rm LLM}(\mathcal{P},\mathcal{T}_{i},\{\mathcal{T}_{j}|v_{j}\in\mathcal{N}_{(u_{i})}\}).$
@@
2.2. An Overview of the Proposed Framework
@@
Figure 2. The overall framework of the proposed TokenRec, which consists of the masked vector-quantized tokenizer with a $K$-way encoder for item ID tokenization, and generative retrieval for recommendation. Note that we detail the item MQ-Tokenizer while omitting the user MQ-Tokenizer for simplicity.
2.3. Masked Vector-Quantized Tokenizers for Users and Items
Instead of assigning each user and item with a specific token (causing a significant increase in vocabulary size), we propose a novel strategy to tokenize users and items to align with natural language, where quantization techniques are developed to represent each user and item with certain discrete indices (i.e., tokens). More specifically, as one of the most representative quantization techniques, Vector Quantization (VQ) aims to convert inputs into a set of discrete tokens (i.e., codes) and learn a discrete latent representation space (i.e., codebook) by reconstructing the original inputs (Esser et al., 2021; Ramesh et al., 2021; Van Den Oord et al., 2017). Moreover, in order to capture high-order collaborative knowledge from users’ interaction history, we propose to conduct the vector quantization on the well-trained representations learned from advanced GNN-based recommendations. However, naively applying vector quantization to index/tokenize users and items is fraught with critical challenges due to poor generalization capability. For example, certain noise and propagation errors in the cascade process of user and item tokenizations in the LLMs greatly hinder the expressiveness of the vanilla VQ.
@@ 2.3. Masked Vector-Quantized Tokenizers for Users and Items
2.3.1. Collaborative Knowledge
The primary objective of our proposed MQ-Tokenizer lies in capturing high-order collaborative knowledge into latent representations through vector quantization. Collaborative knowledge has been widely proven informative for predicting user preferences in recommender systems (Wang et al., 2019; Fan et al., 2019b), as it reveals the in-depth behavioral similarity among users (or items). However, it is challenging for LLMs to map textual descriptions of the high-order collaborative relationships between users and items explicitly with natural language (Wang et al., 2024a). Meanwhile, as the most advanced collaborative filtering techniques, graph neural networks (GNNs) are proposed to learn user and item representations by capturing high-order collaborative signals among users and items in recommendations (He et al., 2020; Fan et al., 2022). Despite great potential, it is challenging to fully harness the high-order signals from user-item interactions in natural language for LLM-based recommendations. Thus, we propose to perform vector quantization on the collaborative representations learned by advanced GNN-based methods for indexing/tokenizing users and items. Specifically, our MQ-Tokenizer assigns each user and item a small number of discrete tokens generated by quantizing the GNN-based representations. In other words, it is suggested that users and items in close proximity to each other in the latent space of collaborative knowledge (i.e., user/item representations) are more likely to share similar tokens/indices, which naturally aligns LLMs with recommendations by tokenizing users and items in natural language.
2.3.2. Masking Operation
@@
(3) $\mathbf{p}^{\prime}_{i}=\text{Mask}(\mathbf{p}_{i},\mathcal{E}),\ \mathbf{q}^{\prime}_{j}=\text{Mask}(\mathbf{q}_{j},\mathcal{E}),$
@@ 2.3.3. $K$-way Encoder and Codebook
(4) $\mathbf{a}^{k}_{j}=\text{Enc}^{k}(\mathbf{q}^{\prime}_{j})=\text{MLP}^{k}(\mathbf{q}^{\prime}_{j}),$
@@
(5) $w^{k}_{j}=\text{arg min}_{l}\|\mathbf{a}^{k}_{j}-\mathbf{c}_{l}^{k}\|^{2},$
(6) $\text{Quantize}(\mathbf{a}^{k}_{j})=\mathbf{c}^{k}_{w^{k}_{j}},$
@@ 2.3.4. $K$-to-1 Decoder
(9) $\mathbf{r}_{j}=\text{Dec}(\{w_{j}^{1},w_{j}^{2},...,w_{j}^{K}\})=\text{MLP}(\frac{1}{K}\sum_{k=1}^{K}\mathbf{c}^{k}_{w^{k}_{j}}).$
2.3.5. Learning Objective
@@
(10) $\mathcal{L}^{Item}_{recon}=\|\mathbf{q}_{j}-\mathbf{r}_{j}\|^{2}.$
However, the arg min operation in Eq. (5) is non-differentiable (Berthet et al., 2020; Lorberbom et al., 2019), leading to an intractable computation for the gradient back-propagation in the reconstruction loss $\mathcal{L}^{Item}_{recon}$ during optimization. To this end, a straight-through gradient estimator (Bengio et al., 2013; Van Den Oord et al., 2017) is introduced to directly assign the gradients of the decoder inputs (i.e., the selected tokens’ embeddings) to the encoder outputs (i.e ., the encoded representations), so as to optimize the encoders and decoder.
@@
(11) $\mathcal{L}^{Item}_{cb}$ $=\sum_{k=1}^{K}\|\text{sg}[~{}\text{Enc}^{k}(\mathbf{q}^{\prime}_{j})~{}]-\mathbf{c}^{k}_{w^{k}_{j}}\|^{2},$
@@
(12) $\mathcal{L}^{Item}_{cm}=\sum_{k=1}^{K}\|~{}\text{Enc}^{k}(\mathbf{q}^{\prime}_{j})~{}-\text{sg}[\mathbf{c}^{k}_{w^{k}_{j}}]\|^{2}.$
@@
(13) $\mathcal{L}^{Item}_{MQ}$ $=\mathcal{L}^{Item}_{recon}+\mathcal{L}^{Item}_{cb}+\beta^{Item}\times\mathcal{L}^{Item}_{cm},$
@@
(14) $\mathcal{L}^{User}_{MQ}$ $=\mathcal{L}^{User}_{recon}+\mathcal{L}^{User}_{cb}+\beta^{User}\times\mathcal{L}^{User}_{cm}.$
2.4. Generative Retrieval for Recommendations
@@
2.4.1. Tokenization&Prompts
Tokenization is the most crucial process of splitting the textual input and output into smaller units that can be processed by language models. In general, LLMs have thousands of tokens in their vocabularies (i.e., in-vocabulary words). For instance, a representative LLM, LLaMA (Touvron et al., 2023), has a vocabulary size of 32,000. Nonetheless, the number of tokens in LLMs remains relatively small compared to the enormous number of users and items in real-world recommendation scenarios, often reaching millions or billions. Therefore, we introduce out-of-vocabulary tokens learned from the proposed user and item MQ-Tokenizers to facilitate the tokenization of user and item IDs. Using our proposed MQ-Tokenizers, only a limited number of $K\times L$ out-of-vocabulary (OOV) tokens are required to tokenize millions or billions of users or items effectively. For example, in our experiments (see Section 3), we can use only 1,536 (i.e., 3 $\times$ 512) OOV tokens to tokenize a total of 39,387 items in the Amazon-Clothing dataset. The vocabulary expansion for tokenizing users and items by MQ-Tokenizers is much more efficient and affordable for LLM-based recommendations. In other words, textual input contents are tokenized by the LLM tokenizer (e.g., SentencePiece), while user and item IDs are tokenized into $K$ discrete tokens (i.e., out-of-vocabulary tokens in LLMs) via the designed user and item MQ-Tokenizers, respectively.
To further enhance the capabilities of LLMs, prompting has been developed to provide explicit guidance to LLMs, enabling them to comprehend textual inputs better and make predictions for downstream tasks (Zhao et al., 2023, 2024). Hence, to generalize to unseen prompts, we design several prompts instructing the LLM backbone to understand users’ preferences for making recommendations. In these prompts, the ID tokens provided by the well-established MQ-Tokenizers are used to represent users and items in the language space of LLM-based recommender systems. For example, given the sub-codebook number $K=3$, two representative prompts in our method and their user&item tokenizations are defined as follows:

### v2

@@ Abstract
There is a growing interest in utilizing large language models (LLMs) to advance next-generation Recommender Systems (RecSys), driven by their outstanding language understanding and reasoning capabilities. In this scenario, tokenizing users and items becomes essential for ensuring seamless alignment of LLMs with recommendations. While studies have made progress in representing users and items using textual contents or latent representations, challenges remain in capturing high-order collaborative knowledge into discrete tokens compatible with LLMs and generalizing to unseen users/items. To address these challenges, we propose a novel framework called TokenRec, which introduces an effective ID tokenization strategy and an efficient retrieval paradigm for LLM-based recommendations. Our tokenization strategy involves quantizing the masked user/item representations learned from collaborative filtering into discrete tokens, thus achieving smooth incorporation of high-order collaborative knowledge and generalizable tokenization of users and items for LLM-based RecSys. Meanwhile, our generative retrieval paradigm is designed to efficiently recommend top-K items for users, eliminating the need for the time-consuming auto-regressive decoding and beam search processes used by LLMs, thus significantly reducing inference time. Comprehensive experiments validate the effectiveness of the proposed methods, demonstrating that TokenRec outperforms competitive benchmarks, including both traditional recommender systems and emerging LLM-based recommender systems.
Index Terms:
@@
Recommender Systems, Large Language Models, ID Tokenization, Vector Quantization, Collaborative Filtering.
I Introduction
@@
As a prominent branch in the data mining field, recommender systems (RecSys) serve as an indispensable and effective technique in addressing information overload problems and enriching user experience across diverse applications [1, 2], such as e-commerce, job search, and social media platforms. To provide personalized recommendations that accord with user preferences, one of the most representative techniques is collaborative filtering (CF), which aims to capture collaborative knowledge by modeling the history of user-item interactions [3]. For example, as the most classic CF method, Matrix Factorization (MF) [4] decomposes the use-item matrix into two low-rank matrices (i.e., user&item representations) to represent each user and item with a unique ID (namely ID indexing/tokenizing), and calculates matching scores to predict user behaviors and generate recommendations via the inner product between their representations. Due to the superior ability in representation learning on graph-structured data, Graph Neural Networks (GNNs) (e.g., LightGCN [5] and GTN [6]) have been recently employed to significantly enhance CF by capturing higher-order collaborative knowledge on user-item interaction graphs for recommendations in IDs manners. The main idea of the existing methods is to obtain learnable representations (i.e., token embeddings) for discrete user&item IDs (i.e., tokens).
@@
Figure 1: Comparison of ID tokenization methods in LLM-based recommendations. Unlike the existing methods, our approach can tokenize users and items with LLM-compatible tokens by leveraging high-order collaborative knowledge.
@@
More recently, the rapid development of Large Language Models (LLMs) techniques (e.g., ChatGPT and LLaMA) has showcased notable milestones for revolutionizing natural language processing techniques [7, 8]. Technically, the unprecedented power of LLMs can be attributed to the scaling up of model/parameter size alongside a tremendous amount of training corpus. In particular, LLMs equipped with billion-scale parameters have exhibited unprecedented language understanding and generation abilities, along with remarkable generalization capability and in-context learning skills that facilitate LLMs to better generalize to unseen tasks and domains [8]. Given the emerging trends and aforementioned advancements of LLMs, LLM-empowered recommender systems have drawn increasing attention from recent studies and demonstrated distinctive abilities for advancing recommender systems [2]. For example, ChatRec [9] integrates ChatGPT into conversational recommendations, which enables users to deliver their explicit preferences in natural language. P5 [10] introduces an LLM-based recommendation framework that unifies diverse recommendation tasks by multi-task prompt-based pre-training.
@@
Unlike the ID-based recommendation methods (e.g., MF and GNN-based), users&items tokenization is one of the most critical steps to take advantage of LLMs within recommendations, as shown in Figure 1. To be more specific, the naive approach known as Independent Indexing (IID) assigns special tokens (i.e., ID) to tokenize each user and item within language models directly. However, this approach becomes infeasible and unrealistic when dealing with large-scale real-world recommender systems, in which the sizes of users and items typically reach billions, dramatically expanding the token vocabulary in LLMs. Meanwhile, as a natural solution, textual title indexing is proposed to utilize LLMs’ in-vocabulary tokens to represent items based on their titles and descriptions, such as the example ”Apple iPhone 15, 256 GB, black”, thus avoiding vocabulary size explosion in LLMs [9, 11]. To achieve a closer alignment between recommendations and natural language, P5 employs whole-word embedding [12] to indicate whether consecutive sub-word tokens originate from the same entities (i.e., user/item). Some studies use continuous embedding (i.e., soft indexing) learned from encoders to represent users&items in LLM-based recommendations [13, 14]. Despite the success mentioned above, the majority of existing methods for users&items tokenization for LLM-based recommender systems still have several limitations. For example, the use of whole-word embedding cannot effectively capture high-order collaborative knowledge and generalize well to unseen users/items for recommendation. In addition, due to the nature of discrete tokens in language models, using continuous indexing makes it challenging to align LLMs in recommender systems closely.
@@
The remainder of this paper is structured as follows. Section II introduces the proposed approach, which is evaluated in Section III. Then, Section IV summarizes the recent development of collaborative filtering and LLM-based RecSys. Finally, conclusions are drawn in Section V.
II The Proposed Method
@@ II-A Notations and Definitions
$\mathbf{z}_{i}={\rm LLM}(\mathcal{P},\mathcal{T}_{i},\{\mathcal{T}_{j}|v_{j}\in\mathcal{N}_{(u_{i})}\}).$ (1)
@@
II-B An Overview of the Proposed Framework
@@
Figure 2: The overall framework of the proposed TokenRec, which consists of the masked vector-quantized tokenizer with a $K$-way encoder for item ID tokenization and the generative retrieval paradigm for recommendation generation. Note that we detail the item MQ-Tokenizer while omitting the user MQ-Tokenizer for simplicity.
II-C Masked Vector-Quantized Tokenizers for Users and Items
Instead of assigning each user and item with a specific token (causing a significant increase in vocabulary size), we propose a novel strategy to tokenize users and items to align with natural language, where quantization techniques are developed to represent each user and item with certain discrete indices (i.e., tokens). More specifically, as one of the most representative quantization techniques, Vector Quantization (VQ) aims to convert inputs into a set of discrete tokens (i.e., codes) and learn a discrete latent representation space (i.e., codebook) by reconstructing the original inputs [15, 16, 17]. Moreover, in order to capture high-order collaborative knowledge from users’ interaction history, we propose to conduct the vector quantization on the well-trained representations learned from advanced GNN-based recommendations. However, naively applying vector quantization to index/tokenize users and items is fraught with critical challenges due to poor generalization capability. For example, certain noise and propagation errors in the cascade process of user and item tokenizations in the LLMs greatly hinder the expressiveness of the vanilla VQ.
@@ II-C Masked Vector-Quantized Tokenizers for Users and Items
II-C1 Collaborative Knowledge
The primary objective of our proposed MQ-Tokenizer lies in capturing high-order collaborative knowledge into latent representations through vector quantization. Collaborative knowledge has been widely proven informative for predicting user preferences in recommender systems [18, 3], as it reveals the in-depth behavioral similarity among users (or items). However, it is challenging for LLMs to map textual descriptions of the high-order collaborative relationships between users and items explicitly with natural language [19]. Meanwhile, as the most advanced collaborative filtering techniques, graph neural networks (GNNs) are proposed to learn user and item representations by capturing high-order collaborative signals among users and items in recommendations [5, 6]. Despite great potential, it is challenging to fully harness the high-order signals from user-item interactions in natural language for LLM-based recommendations. Thus, we propose to perform vector quantization on the collaborative representations learned by advanced GNN-based methods for indexing/tokenizing users and items. Specifically, our MQ-Tokenizer assigns each user and item a small number of discrete tokens generated by quantizing the GNN-based representations. In other words, it is suggested that users and items in close proximity to each other in the latent space of collaborative knowledge (i.e., user/item representations) are more likely to share similar tokens/indices, which naturally aligns LLMs with recommendations by tokenizing users and items in natural language.
II-C2 Masking Operation
@@
$\mathbf{p}^{\prime}_{i}=\text{Mask}(\mathbf{p}_{i},\mathcal{E}),\ \mathbf{q}^{\prime}_{j}=\text{Mask}(\mathbf{q}_{j},\mathcal{E}),$ (3)
@@ II-C3 $K$-way Encoder and Codebook
$\mathbf{a}^{k}_{j}=\text{Enc}^{k}(\mathbf{q}^{\prime}_{j})=\text{MLP}^{k}(\mathbf{q}^{\prime}_{j}),$ (4)
@@
$w^{k}_{j}=\text{arg min}_{l}\|\mathbf{a}^{k}_{j}-\mathbf{c}_{l}^{k}\|^{2},$ (5)
$\text{Quantize}(\mathbf{a}^{k}_{j})=\mathbf{c}^{k}_{w^{k}_{j}},$ (6)
@@ II-C4 $K$-to-1 Decoder
$\mathbf{r}_{j}=\text{Dec}(\{w_{j}^{1},w_{j}^{2},...,w_{j}^{K}\})=\text{MLP}(\frac{1}{K}\sum_{k=1}^{K}\mathbf{c}^{k}_{w^{k}_{j}}).$ (9)
II-C5 Learning Objective
@@
$\mathcal{L}^{Item}_{recon}=\|\mathbf{q}_{j}-\mathbf{r}_{j}\|^{2}.$ (10)
However, the arg min operation in Eq. (5) is non-differentiable [20, 21], leading to an intractable computation for the gradient back-propagation in the reconstruction loss $\mathcal{L}^{Item}_{recon}$ during optimization. To this end, a straight-through gradient estimator [22, 17] is introduced to directly assign the gradients of the decoder inputs (i.e., the selected tokens’ embeddings) to the encoder outputs (i.e ., the encoded representations), so as to optimize the encoders and decoder.
@@
$\mathcal{L}^{Item}_{cb}$ $=\sum_{k=1}^{K}\|\text{sg}[~{}\text{Enc}^{k}(\mathbf{q}^{\prime}_{j})~{}]-\mathbf{c}^{k}_{w^{k}_{j}}\|^{2},$ (11)
@@
$\mathcal{L}^{Item}_{cm}=\sum_{k=1}^{K}\|~{}\text{Enc}^{k}(\mathbf{q}^{\prime}_{j})~{}-\text{sg}[\mathbf{c}^{k}_{w^{k}_{j}}]\|^{2}.$ (12)
@@
$\mathcal{L}^{Item}_{MQ}$ $=\mathcal{L}^{Item}_{recon}+\mathcal{L}^{Item}_{cb}+\beta^{Item}\times\mathcal{L}^{Item}_{cm},$ (13)
@@
$\mathcal{L}^{User}_{MQ}$ $=\mathcal{L}^{User}_{recon}+\mathcal{L}^{User}_{cb}+\beta^{User}\times\mathcal{L}^{User}_{cm}.$ (14)
II-D Generative Retrieval for Recommendations
@@
II-D1 Tokenization&Prompts
Tokenization is the most crucial process of splitting the textual input and output into smaller units that language models can process. In general, LLMs have thousands of tokens in their vocabularies (i.e., in-vocabulary words). For instance, a representative LLM, LLaMA [23], has a vocabulary size of 32,000. Nonetheless, the number of tokens in LLMs remains relatively small compared to the enormous number of users and items in real-world recommendation scenarios, often reaching millions or billions. Therefore, we introduce out-of-vocabulary tokens learned from the proposed user and item MQ-Tokenizers to facilitate the tokenization of user and item IDs. Using the MQ-Tokenizers, only a limited number of $K\times L$ out-of-vocabulary (OOV) tokens are required to tokenize millions or billions of users or items effectively. For example, in our experiments (see Section III), we can use only 1,536 (i.e., 3 $\times$ 512) OOV tokens to tokenize a total of 39,387 items in the Amazon-Clothing dataset. The vocabulary expansion for tokenizing users and items by MQ-Tokenizers is much more efficient and affordable for LLM-based recommendations. In other words, textual contents are tokenized by the LLM tokenizer (e.g., SentencePiece), while user and item IDs are tokenized into $K$ discrete OOV tokens via the corresponding MQ-Tokenizers, respectively.
To further enhance the capabilities of LLMs, prompting has been developed to provide explicit guidance to LLMs, enabling them to make better predictions for downstream tasks [8, 2]. Building upon this insight, we design several prompts instructing the LLM backbone to understand users’ preferences for making recommendations. In these prompts, the ID tokens provided by the well-established MQ-Tokenizers are used to represent users and items in the language space of LLM-based recommender systems. For example, given the sub-codebook number $K=3$, two representative prompts in our method and their user&item tokenizations are defined as follows: